OpenStudy (anonymous):

100/50 write the prime fractorization of the numerator and the denominator. then write the fraction in lowest terms

OpenStudy (tpaulus):

100/50 = 50x2/50 = 2

OpenStudy (anonymous):

thanks, but how did u get that answer

OpenStudy (tpaulus):

100 = 50x2, and 50/50 is one, so 50/50 x2 = 2

OpenStudy (anonymous):

\[ 100 = 2\times 50 = 2\times 2\times 25 = 2\times 2\times 5\times 5 = 2^25^2 \]Then \[ 50 = 2\times 25 = 2\times 5\times 5 = 2^15^2 \]

OpenStudy (anonymous):

So \[ \frac{100}{50} =\frac{2^2\cancel{5^2}}{2^1\cancel{5^{2}}} = \frac{2^{\cancel 2 1}}{\cancel{2^1}} = 2 \]

OpenStudy (anonymous):

thks woi

OpenStudy (anonymous):

wio

OpenStudy (anonymous):

Or just divide